A sample of hydrogen gas has a volume of 7.49 L at a pressure of 22.0 atm and a temperature of 32.0°C. What volume of gaseous water is produced by the following reaction at 125.0°C and 0.975 atm, if all the hydrogen gas reacts with iron (III) oxide? 

1
Expert's answer
2021-10-30T11:16:07-0400
  • The volume of hydrogen gas is 7.49 L.
  • The pressure of hydrogen gas is 22.0 atm.
  • The temperature of hydrogen gas is 32.0"\u00b0c"
  • 305.15
  • K
  • 32.0∘C(32.0+273.15)K=305.15K
  • .
  • The pressure of gaseous water is 0.975 atm.
  • The temperature of gaseous water is 125.0
  • C
  • (
  • 125.0
  • +
  • 273.15
  • )
  • K
  • =
  • 398.15
  • K
  • 125.0∘C(125.0+273.15)K=398.15K
  • .


The balanced chemical equation is written as given below:

F

e

2

O

3

(

s

)

+

3

H

2

(

g

)

2

F

e

(

s

)

+

3

H

2

O

(

g

)

Fe2O3(s)+3H2(g)→2Fe(s)+3H2O(g)

The number of moles of hydrogen gas is calculated by using the ideal gas equation as given below:

P

V

=

n

R

T

PV=nRT

Where,

  • P is the pressure of a gas, i.e., 22.0 atm.
  • T is the temperature of a gas, i.e., 305.15 K.
  • V is the volume of a gas, i.e.,7.49 L.
  • n is the number of moles of a gas.
  • R is the universal gas constant (
  • R
  • =
  • 0.0821
  • L
  • a
  • t
  • m
  • K
  • 1
  • m
  • o
  • l
  • 1
  • )
  • (R=0.0821L⋅atm⋅K−1⋅mol−1)
  • .


On substituting the values in the ideal gas equation, we get

P

V

=

n

R

T

22.0

a

t

m

×

7.49

L

=

n

×

0.0821

L

a

t

m

K

1

m

o

l

1

×

305.15

K

PV=nRT22.0atm×7.49L=n×0.0821L⋅atm⋅K−1⋅mol−1×305.15K

On rearranging and simplifying the ideal gas equation, we get

n

H

2

=

22.0

a

t

m

×

7.49

L


0.0821

L

a

t

m

K

1

m

o

l

1

×

305.15

K

=

164.8


25.05

=

6.58

m

o

l

nH2=22.0atm×7.49L0.0821L⋅atm⋅K−1⋅mol−1×305.15K=164.825.05=6.58mol


According to stoichiometry, 3 mol of hydrogen gas produces 3 mol of water. Therefore, the number of moles of water that can be produced from 6.58 mol of hydrogen gas is calculated as given below:

M

o

l

e

s

o

f

H

2

O

=

6.58

m

o

l

o

f

H

2

×

3

m

o

l

H

2

O


3

m

o

l

H

2

=

6.58

m

o

l

MolesofH2O=6.58molofH2×3molH2O3molH2=6.58mol


Hence the volume of water is calculated by using the formula as given below:

P

V

=

n

R

T

PV=nRT

Where,

  • P is the pressure of a gas, i.e., 0.975 atm.
  • T is the temperature of a gas, i.e., 398.15 K.
  • V is the volume of a gas.
  • n is the number of moles of a gas, i.e., 6.58 mol.
  • R is the universal gas constant (
  • R
  • =
  • 0.0821
  • L
  • a
  • t
  • m
  • K
  • 1
  • m
  • o
  • l
  • 1
  • )
  • (R=0.0821L⋅atm⋅K−1⋅mol−1)
  • .


On substituting the values in the ideal gas equation, we get

P

V

=

n

R

T

0.975

a

t

m

×

V

H

2

O

=

6.58

m

o

l

×

0.0821

L

a

t

m

K

1

m

o

l

1

×

398.15

K

PV=nRT0.975atm×VH2O=6.58mol×0.0821L⋅atm⋅K−1⋅mol−1×398.15K

On rearranging and simplifying the ideal gas equation, we get

V

H

2

O

=

6.58

m

o

l

×

0.0821

L

a

t

m

K

1

m

o

l

1

×

398.15

K


0.975

a

t

m

=

215.1


0.975

=

220.6

L

221

L

VH2O=6.58mol×0.0821L⋅atm⋅K−1⋅mol−1×398.15K0.975atm=215.10.975=220.6L≈221L


Thus, the volume of gaseous water is 221 L.

.